OpenStudy (anonymous):

An observer (O) spots a plane (P) taking off from a local airport and flying at a 33° angle horizontal to her line of sight and located directly above a tower (T). The observer also notices a bird (B) circling directly above her. If the distance from the plane(P) to the tower (T) is 7,000 ft., how far is the bird (B) from the plane (P)? 3815 feet 5873 feet 8343 feet 10,779 feet

OpenStudy (anonymous):

@Hero can you help

OpenStudy (aum):



OpenStudy (aum):

Find distance BP. BP = OT tan(33) = PT / OT. OT = PT / tan(33) = 7000 / tan(33) = ?

OpenStudy (anonymous):

10779.1

OpenStudy (anonymous):

thank you

OpenStudy (aum):

you are welcome.